REU 2005: SCHEDULE FOR WEEKS THREE AND FOUR

Morning talks: Ryerson Annex 276 Afternoon talks: Eckhart 206

Note the changed times of the afternoon talks Tuesday, July 5: YSP needs the 4:00 – 5:00 hour for analysis of diagnostic tests.

Date  YSP  Abert/Babai  Abert  Babai  Bader  Gordon
Apprentice  Problems Discrete
Ryer. Annex 276 Eck 206 Eck 206 Eck 206 Eck 206
Tue, July 5 9:30–2:30  9:30–12:00  2:45–4:00  5:00–6:00
Wed, July 6 9:30–2:30  9:30–12:00  2:45–4:15  4:30–6:00
Thu, July 7 9:30–2:30  9:30–12:00  2:45–4:15  4:30–6:00
Fri, July 8 9:30–2:30  9:30–12:00  2:45–4:15  4:30–6:00
Date  YSP  Abert/Babai  Abert  Babai Muchnik Gordon
Apprentice  Problems Discrete
Ryer. Annex 276 Eck 206 Eck 206 Eck 206 Eck 206
Mon, July 11 9:30–2:30  9:30–12:00  2:45–4:15  4:30–6:00
Tue, July 12 9:30–2:30  9:30–12:00  2:45–4:15  4:30–6:00
Wed, July 13 9:30–2:30  9:30–12:00  2:45–4:15  4:30–6:00
Thu, July 14 9:30–2:30  9:30–12:00  2:45–4:15  4:30–6:00
Fri, July 15 9:30–2:30  9:30–12:00  4:30–6:00 2:45–4:15

Here is the schedule in summary form:

THIRD WEEK:

YSP 9:30–2:30 TWThF Abert/Babai (Apprentice) 9:30–12:00 TWThF Abert (Problems) 2:45–4:00 T, 2:45–4:15 Th Babai (Discrete mathematics) 2:45–4:15 WF Bader (Hyperbolic manifolds) 5:00–6:00 T, 4:30–6:00 Th Gordon (Nonlinear waves) 4:30–6:00 WF

FOURTH WEEK:

YSP 9:30–2:30 M–F Abert/Babai (Apprentice) 9:30–12:00 M–F Abert (Problems) 2:45–4:15 TTh Babai (Discrete mathematics) 2:45–4:15 MW Muchnik (Hyperbolic manifolds) 4:30–6:00 TThF Gordon (Nonlinear waves) 4:30–6:00 MW and 2:45–4:15 F